# Seeded Poisson Factorization: Leveraging domain knowledge to fit topic models

### v1

@@ 4.3 Robustness checks
In Scenarios 3, 4, 6 and 7 the learning rate is reduced to 0.01 from 0.1 in the base scenario. Results indicate that in particular lowering the learning rate increases the number of epochs required for the negative $\mathsf{ELBO}$ to converge, suggesting slower optimization (see Scenario 4).
@@
Dropping the seed words for one topic but otherwise keeping the number of topics (Scenario 8) results in them same good performance than in the base scenario. In the case where ‘Grocery’ topic is excluded and an unseeded topic is added, SPF allocates 4 82648264\,8264 826 customer reviews to the unseeded category, accurately identifying 2 33123312\,3312 331 out of 2 47124712\,4712 471 instances as belonging to the ‘Grocery’ category. Inspecting the words with the highest topic-term intensities for the unseeded topic indicates that this topic effectively captures the ‘Grocery’ topic. High-intensity words emerging in this setting are ‘taste’ (12.84), ‘tea’ (12.13), and ‘flavor’ (8.78).
In Scenario 9, adding an additional unseeded topic leads to reduced model performance ($0.67$ compared to $0.73$ in the baseline scenario), as the fixed number of six product categories means that assigning a customer review to the unseeded topic constitutes a mis-classification in this context. However, an analysis of the topic-term distributions for the unseeded topic reveals that SPF assigns reviews to the unseeded topic when customers primarily discuss the purchasing process rather than specific product characteristics. To give an example, terms such as ‘time’, ‘shipping’, ‘store’, ‘order’ and ‘online’ exhibit high intensities within the unseeded topic. SPF was hence capable to identify the additional latent topic present in customer reviews which relates to the purchasing and delivery experiences, which, while not tied to specific product categories, is also of significant business relevance. Overall these findings illustrate SPF’s robust capability to discern meaningful latent topics even in the absence of comprehensive domain-specific seeding.
@@
[…]
@@ 5 Discussion
Traditional topic models often struggle to align the latent topics they derive with pre-specified concepts of interest (see, e.g., Eshima et al.,, 2024). To address these limitations, we extend the Poisson factorization topic model with a seeded approach. The seeded approach guides the inference of topics, avoiding the need for manual labeling, but also enables the use of topic models for text classification where labeled text data are not available but the classes for categorization are readily characterized by a set of relevant words. Seeding modifies the prior distribution of the topic-term distributions by assigning higher rates a-priori to the relevant words associated with their respective topics.
Our empirical findings demonstrate that integrating domain knowledge into the model specification significantly enhances the capability of topic models to extract meaningful topic-term intensities, thereby improving the interpretability and understanding of topics. Additionally, by applying a Naive Bayes classifier based on the fitted document-topic distributions, we are able to classify documents in an automatic way. Experiments on a pre-labeled dataset reveal that the SPF topic model not only improves topic interpretability but also achieves superior predictive performance compared to alternative seeded topic modeling approaches. By combining the computational efficiency of VI techniques with the prior knowledge of domain experts in a PF framework, SPF enables a robust and effective system for document classification. This synergy improves the overall quality and utility of the classification process, making it more reliable and actionable for a wide range of applications.
Acknowledgments

### v2

@@ 4.1.4 Robustness checks
In Scenarios 3, 4, 6 and 7 the learning rate is reduced to 0.01 from 0.1 in the base scenario. The results indicate that in particular lowering the learning rate increases the number of epochs required for the negative $\mathsf{ELBO}$ to converge, suggesting slower optimization (see Scenario 4).
@@
Dropping the seed words for one topic but otherwise keeping the number of topics (Scenario 8) results in the same good performance as in the base scenario. In the case where the ‘Grocery’ topic is excluded and an unseeded topic is added, SPF allocates 4 8264\,826 customer reviews to the unseeded category, accurately identifying 2 3312\,331 out of 2 4712\,471 instances as belonging to the ‘Grocery’ category. Inspecting the words with the highest topic-term intensities for the unseeded topic indicates that this topic effectively captures the ‘Grocery’ topic. High-intensity words emerging in this setting are ‘taste’ (12.84), ‘tea’ (12.13), and ‘flavor’ (8.78).
In Scenario 9, adding an additional unseeded topic leads to reduced model performance ($0.67$ compared to $0.73$ in the baseline scenario), as the fixed number of six product categories means that assigning a customer review to the unseeded topic constitutes a misclassification in this context. However, an analysis of topic-term distributions for the unseeded topic reveals that SPF assigns reviews to the unseeded topic when customers primarily discuss the purchasing process rather than specific product characteristics. To give an example, terms such as ‘time’, ‘shipping’, ‘store’, ‘order’ and ‘online’ exhibit high intensities within the unseeded topic. SPF was therefore capable to identify the additional latent topic present in customer reviews which relates to purchasing and delivery experiences, which – while not tied to specific product categories – is also of significant business relevance. Overall these findings illustrate SPF’s robust capability to discern meaningful latent topics even in the absence of comprehensive domain-specific seeding.
@@
We explored the impact of varying the a-priori relevance assigned to seed words by adjusting the shape parameter $c$ of the prior of the seeded topic-term intensities (Scenarios 10–13). Our findings indicate that the selection of $c$ has minimal impact when the data size is large, i.e., D=30  [^0] D=30\,000. In this case, the influence of the prior is outweighed by the substantial information in the data (see Scenarios 10 and 11 with an accuracy of 0.73 each).
Changing the value of the shape parameter $c$ for a smaller dataset (D=1  [^0] D=1\,000) indicates that this has some effect on model accuracy. In this case, the model accuracy is slightly higher for a more informative prior setting compared to a setting where only a small amount of additional weight is imposed on the seed words, i.e., an accuracy of $0.64$ is obtained in Scenario 13 compared to $0.62$ in Scenario 12. This observation underscores the importance of balancing prior informativeness with dataset size for optimal model performance.
4.2 Application to four benchmark datasets
We evaluate SPF model’s performance and general applicability using four additional publicly available text corpora. These corpora encompass various domains and in particular also feature a wide range of different number of topics to be estimated, which allows to further validate the SPF approach. Pre-processing involves normalization (conversion to lowercase), tokenization and the removal of documents that result in zero tokens after processing. Dataset statistics are summarized in Table 7.
  • Banking [71] [^3] https://huggingface.co/datasets/banking77 is a dataset containing customer service queries from the banking sector. It includes over 10 00010\,000 labeled queries across 77 fine-grained banking-related intents. We use the query text as the document input and align the pre-defined intent labels with the seed topics. Stop words are retained during pre-processing, resulting in a vocabulary size of V=2 320V=2\,320.
  • DBPedia [118] [^4] https://huggingface.co/datasets/dbpedia_14 provides structured information extracted from Wikipedia. For our experiments, we use a classified subset consisting of 14 non-overlapping categories (e.g., company, artist, place). The title and abstract fields are concatenated to form the document text, and the 14 class labels are used to derive the seed words. During pre-processing, stop words are removed, and the vocabulary is limited to the top V=25  [^0] V=25\,000 most frequent terms.
  • 20NG [93] [^5] https://scikit-learn.org/0.19/datasets/twenty_newsgroups.html is a well-known dataset composed of posts to 20 different newsgroups. We treat the post content as document and use the newsgroup information (e.g., comp.graphics, sci.space) as classes for the seed words. We also remove stop words during pre-processing and limit the vocabulary to the 25 00025\,000 most frequent terms.
  • Ledgar [111] [^6] https://aclanthology.org/2020.lrec-1.155/ is a large-scale multi-label corpus of legal clauses extracted from SEC filings. It contains almost 100 000 contractual provisions annotated with over 12 000 clause types. For our experiments, we use the LEX-GLUE LEDGAR subset [^7] https://huggingface.co/datasets/coastalcph/lex_glue with around 100 clause types, treating clause texts as documents and derive the seed words based on the clause types. We remove stop words during pre-processing, which results in a vocabulary of 18 47618\,476 unique terms.
Dataset Characteristics Model Algorithm Classification Topic coherence and diversity $D$ # $\varnothing$ K $E$ $\rho$ Acc Prec Rec F1 NPMI UMass $\mathbf{C_{V}}$ Div Banking 10 003 20 10.76 77 SPF 750 0.001 0.72 0.73 0.71 0.71  0.03 -1.97 0.44 0.12 PF 750 0.001 0.04 0.04 0.04 0.04  0.12 -1.70 0.36 0.01 20NG 18 267 5 91.48 20 SPF 550 0.01 0.60 0.62 0.58 0.56  0.02 -1.93 0.70 0.79 PF 550 0.01 0.35 0.36 0.34 0.33 -0.01 -2.02 0.65 0.76 Ledgar 60 000 25 54.43 100 SPF 550 0.0005 0.61 0.52 0.58 0.51  0.05 -2.04 0.71 0.48 PF 550 0.0005 0.02 0.02 0.02 0.02  0.13 -1.39 0.51 0.02 DBPedia 559 975 25 24.74 14 SPF 250 0.01 0.84 0.85 0.84 0.84  0.19 -2.45 0.78 0.95 PF 250 0.01 0.39 0.39 0.39 0.38  0.13 -2.70 0.63 0.76
Table 7: Evaluation of model performance across different datasets. For each dataset, we report key corpus characteristics (# indicates the number of seed terms per topic, $\varnothing$ denotes the average number of words per document), algorithm settings, classification performance (Accuracy – Acc, Precision – Prec, Recall – Rec, F1-score – F1), topic coherence (NPMI, UMass, $C_{V}$) and diversity (Div) for both SPF and PF topic models.
We set the number of topics to the number of categories provided for each dataset (i.e., 77 for Banking, 14 for DBPedia, 20 for 20NG, and 100 for Ledgar). We initialize the hyperparameters of the gamma priors as outlined in Section 2 and train both models with learning rates and number of epochs, specifically selected for each corpus, closely monitoring the convergence of the ELBO. The settings used and results obtained are reported in Table 7. The experiments are run on the hardware described in Section 3.3. For coherence and topic diversity evaluation, we use the top-10 ranked words per topic.
SPF achieves strong predictive performance across all datasets, with the accuracy scores varying between 0.60 and 0.84. A particularly high accuracy score is obtained for the large-scale DBPedia dataset ($0.84$), which has 14 categories. Also the accuracy score of $0.72$ obtained for the Banking dataset is impressive, in particular given the 77 categories to which one assigns. This excellent classification performance indicates that SPF maintains competitive results, demonstrating robustness to variations in corpus size, document length, number of topics, and number of seed terms. In contrast, the PF model, which is completely unsupervised, results consistently in lower classification performance, with accuracy scores ranging from $0.02$ on Ledgar to $0.39$ on DBPedia. This highlights the substantial benefit of incorporating minimal domain knowledge through seed terms, as in SPF.
The topic coherence and diversity scores reveal an expected trade-off. For example, DBPedia, which yields the best classification performance for SPF, also achieves the highest topic diversity ($0.95$) and $C_{V}$ coherence ($0.78$), while scoring poorly on UMass (2.45-2.45). PF, on the other hand, tends to produce slightly higher NPMI and UMass scores in some settings (e.g., Banking), likely due to its unsupervised optimization of topic structure rather than predictive alignment. However, this comes at the cost of classification performance. Overall, while PF occasionally produces more coherent or diverse topic structures according to select metrics, SPF clearly outperforms it in classification tasks. The results confirm that even limited supervision, as provided by seed terms, significantly enhances predictive accuracy with only a modest impact on topic quality.
@@ 5 Discussion
Traditional topic models often struggle to align the latent topics they derive with pre-specified concepts of interest [76, see, e.g.,]. To address these limitations, we extend the PF topic model with a seeded approach. The seeded approach guides the inference of topics, avoiding the need for manual labeling, but also enables the use of topic models for text classification where labeled text data are not available but the classes for categorization are readily characterized by a set of relevant words. Seeding modifies the prior distribution of the topic-term distributions by assigning higher rates a-priori to the relevant words associated with their respective topics.
Our empirical findings demonstrate that integrating domain knowledge into the model specification significantly enhances the capability of topic models to extract meaningful topic-term intensities, thereby improving the understanding of topics. Additionally, by applying a Naive Bayes classifier based on the fitted document-topic distributions, we are able to classify documents automatically. Experiments on datasets with known categorizations reveal that the SPF topic model achieves superior classification performance compared to alternative seeded probabilistic topic modeling approaches. By combining the computational efficiency of VI techniques with the prior knowledge of domain experts in a PF framework, SPF enables a robust and effective system for document classification. This synergy improves the overall quality and utility of the classification process, making it more reliable and actionable for a wide range of applications.
SPF relies on the bag-of-words assumption to allow straightforward inclusion of domain knowledge and efficient estimation. SPF, however, might benefit, in particular, from recent advances in deep learning methods, including transformer or Mamba models, by exploiting their capabilities for the improved derivation of sets of terms to be used as seed words, see [99] and [120].
Acknowledgements